Find the angle between the lines 3x+y+12=0 and x+2y-1=0 and the equations of the line drawn perpendicular to them from the point (3,-2)

Grade:10

1 Answers

Vikas TU
4 years ago
Dear student 
slope of line  = -A/B 
m1 = -3 m2 = -0.5 
TanQ  = m1m2/1-m1m2
Put the value of m1 and m2 
Q = tan inverse (-3)
Sope oof line perpendicular to line 1 = -1/3 
Slope of line perpendicular to line 2 = -2 
Equation 1 = y+2 = -1/3 (x-3)
Equation 2 = y+2 = -2(x-3)